OpenStudy (anonymous):

multiply (3 sqrt 5+1)(2 sqrt 5-3)

OpenStudy (anonymous):

like this? \[(3\sqrt{5})/1 \times (2\sqrt{5})-3\]

OpenStudy (anonymous):

or is it another grouping?

OpenStudy (anonymous):

\[(3\sqrt{5} + 1)(2\sqrt{5} - 3) \implies 3\sqrt{5}(2\sqrt{5} - 3) + 1(2\sqrt{5} - 3) \implies?\]

OpenStudy (anonymous):

(35√+1)(25√−3)<-- its this one

OpenStudy (anonymous):

Is it \[(35\sqrt{1})(25\sqrt{-3})\]

OpenStudy (anonymous):

no its (3√5+1)(2√5−3)

OpenStudy (anonymous):

Then my drawing above is correct?

OpenStudy (anonymous):

I mean my first post.

OpenStudy (anonymous):

its not 35 and 25 ...

OpenStudy (anonymous):

My first post...not my second...

OpenStudy (anonymous):

nope ... its +1 not divided by 1 . this is the way you write the equation (3√5+1)(2√5−3)

OpenStudy (anonymous):

MY first post...not @hellotiff 's, MY first post. THis one: \[(3\sqrt{5} + 1)(2\sqrt{5} - 3) \implies 3\sqrt{5}(2\sqrt{5} - 3) + 1(2\sqrt{5} - 3) \implies?\]

OpenStudy (anonymous):

oh goodness i didnt notice it waqs someone else lmao yes !

OpenStudy (anonymous):

Alright. \[3\sqrt{5}(2\sqrt{5} - 3) + 1(2\sqrt{5} - 3) \implies 30 - 9\sqrt{5} + 2\sqrt{5} - 3\]Can you simplify?

OpenStudy (anonymous):

27-7√5 ?

OpenStudy (anonymous):

Yup :) Good job!

OpenStudy (anonymous):

thanks ,

OpenStudy (anonymous):

np :)